OpenStudy (anonymous):

A rectangular garden with a width of 42 feet has an area of 2,814 square feet. What is the length of the garden?

OpenStudy (yolomcswagginsggg):

The formula for the area of a rectangle is \(Area = l \times w\).

OpenStudy (yolomcswagginsggg):

Plug in what we know into the formula: \(Area = l \times w\) We have: Width = 42 ft Area = 2814 sqft Plug them into the equation: \(2814 = l \times 42\) Solve for the length: Divide 42 to both sides: \(\dfrac{2814}{42} = \dfrac{l \times 42}{42}\)  The 42's at the right cancel out, leaving you with \(l\) by itself. Now divide 2814/42. \(l = 67\)

OpenStudy (anonymous):

lol